40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1962     Entry time: Mon Oct 30 00:29:09 2017
Author: awade 
Type: DailyProgress 
Category: FSS 
Subject: Issues locking FSS; acromag excitation voltage issues 

I tried to lock both the north and south this afternoon and ran into issues with engaging the lock.  

Two issues. 

First: when I plugged the laser PZT into a HV amp and scanned the voltage I couldn't see an error signal on the oscilloscope for the south path. I didn't get to the bottom of why this is the case.  I got as far as checking the 14.75 MHz modulation was on, checking electronics were powered and connected properly and sticking a high impedance oscilloscope probe directly into the field box.  Tracking down what is going on here is the first order of business tomorrow morning.

Second issue: I also checked the level of the voltage going into the digital switch U1 (pin 20) on the D040105 board in the field box.  This was registering nothing. I tracked this back to no voltage out of the acromag digital logic outputs.  Through various iterations of dissembling the TTFSS Ethernet controller box I found that the culprit was the DC level of the external excitation voltage supplied to the Acromag. It activates on >= 5.1 V but not below 4.9 V supply.  The level supplied by the LM7805 (from 24V line) is below 5.0 V -- many of these regulators are -- so the it  is failing to provide any voltage at the output of the digital outs. I suspect that the plug pack we were previously using was also doing such things intermittently (it looks cheep).  Tomorrow's second order of business is finding a plug pack that supplies a generous 5.1+ volts and connecting barrel connectors into the TTFSS Ethernet module.

If we can't resolve the Ethernet module digital issues we should go ahead and manually lock to take a beat not spectrum tomorrow afternoon.

ELOG V3.1.3-